# Simulation studies on compensation for space-charge-induced half-integer and 3rd-order resonance crossing in HIAF-BRing

### v1

@@ III.1 Half-integer resonance crossing at $192Q_{x}=$ 19
The oscillation of the phase space in Fig. (9) is not as clear as that in FIg. (5). An additional Fig. (10) is introduced to demonstrate the evolution of the action of particles in Fig. (9). For the result with modified RDTs, although its horizontal RMS emittance is almost the same as that without errors, its actions are not. For example, in the middle top subgraph of Fig. (10), the actions of the red particle whose initial $J_{x}$ is $0.8\times 10^{-4}$ m and the orange particle whose initial $J_{x}$ is $0.7\times 10^{-4}$ m decrease while the actions of the green particle whose initial $J_{x}$ is $0.6\times 10^{-4}$ m and the blue particle whose initial $J_{x}$ is $0.5\times 10^{-4}$ m increase. In the middle bottom subgraph, the action of the black halo particle becomes similar to that in the right bottom subgraph after 650 turns, which causes the "slight" $283Q_{y}=$ 28 resonance behavior in the middle bottom subgraph in Fig. (9).
@@
Figure 13: The normalized phase spaces under the tune spread in Fig. (11). The left 2 subgraphs show the phase space without errors. The middle 2 subgraphs show that with modified RDTs. The right 2 subgraphs show that with the scheme of J-PARC-MR.
Figure 14: The action in Fig. (13). The left 2 subgraphs show the phase space without errors. The middle 2 subgraphs show that with modified RDTs. The right 2 subgraphs show that with the scheme of J-PARC-MR.
The Fig. (7-10) could be redrawn into Fig. (11-14) after introducing chromaticity in the simulation. In Fig. (11), the chromaticity introduces an extra tune shift $\xi_{u}\delta$ to each particle, as seen in the left top subgraph in Fig. (6). In Fig. (12), for the result without compensation, the beam loss is rapid, causing the statistical $\varepsilon_{rms,x}$ to decrease. Conversely, the $\varepsilon_{y}$ grows rapidly, resulting in an increase in statistical $\varepsilon_{rms,y}$ despite the beam loss. For the result with raw RDTs, emittance growth could be observed in both horizontal and vertical planes, indicating that the compensation scheme with raw RDTs is not effective for magnetic field errors under space-charge-induced resonance crossing. For the result with modified RDTs, the $\varepsilon_{rms,x}$ is almost the same as that without errors, and the $\varepsilon_{rms,y}$ grows by 2.5%, which is 0.5% larger than growth observed in Fig. (8). In Fig. (13), for the phase space with modified RDTs, the amplitude of the black halo particle becomes even smaller than that without errors. Fig. (15) demonstrates the $J_{u}$ evolution of this particle, showing that there is no action transfer from the vertical to the horizontal plane. This phenomenon is interesting as the mechanism behind this action decrease under the combined effect of space charge, chromaticity, sextupole errors and corresponding compensation is unclear and warrants further study.
@@ IV CONCLUSIONS
In this paper, the key idea is to propose the space charge modification in RDTs, called modified RDTs. The modification is an approximate approach in order to obtain the Twiss of particles under the space-charge-induced resonance crossing, through which the modified RDTs could be obtained. The modified RDTs aim to compensate for the combined effect of space charge and magnetic field imperfections, focusing on particles under resonance crossing. This approach is precise for core particles but imprecise for halo particles.
The feasibility of half-integer and 3rd-order resonance crossing compensated by modified RDTs is confirmed by the simulation, in which the raw RDTs and the J-PARC-MR scheme are also demonstrated for comparison. Among these schemes, the modified RDTs exhibits the best performance. The simulations are conducted both with and without chromaticity. The chromaticity has little influence on the overall compensation with modified RDTs, as evidenced by the almost unchanged RMS emittance evolution.

### v2

@@ III.2 Half-integer resonance crossing at $192Q_{x}=$ 19
The oscillation of the phase space in Fig. (11) is not as clear as that in Fig. (7). An additional Fig. (12) is introduced to demonstrate the evolution of the action of particles in Fig. (11). For the result with modified RDTs, although its horizontal RMS emittance is almost the same as that without errors, its actions are not. For example, in the middle top subgraph of Fig. (12), the actions of the red particle whose initial $J_{x}$ is $0.8\times 10^{-4}$ m and the orange particle whose initial $J_{x}$ is $0.7\times 10^{-4}$ m decrease while the actions of the green particle whose initial $J_{x}$ is $0.6\times 10^{-4}$ m and the blue particle whose initial $J_{x}$ is $0.5\times 10^{-4}$ m increase. In the middle bottom subgraph, the action of the black halo particle becomes similar to that in the right bottom subgraph after 650 turns, which causes the "slight" $283Q_{y}=$ 28 resonance behavior in the middle bottom subgraph in Fig. (11). In the initial 650 turns, this particle does not hit the stop band of $283Q_{y}=$ 28. As the vertical emittance increases, the vertical space charge effect and the tune spread decreases. So this particle hit the stop band and is locked in the resonance islands in the rest of the evolution. For the result with moved RDTs, the stop band is larger than that with modified RDTs. As a result, the black halo particle is locked in the resonance islands in the beginning. Note that particles with the same color in the horizontal or vertical plane are the same particles.
@@
Figure 15: The normalized phase spaces under the tune spread in Fig. (13). The left 2 subgraphs show the phase space without errors. The middle 2 subgraphs show that with modified RDTs. The right 2 subgraphs show that with moved RDTs.
Figure 16: The action in Fig. (15). The left 2 subgraphs show the phase space without errors. The middle 2 subgraphs show that with modified RDTs. The right 2 subgraphs show that with moved RDTs.
The Fig. (9-12) could be redrawn into Fig. (13-16) after introducing chromaticity in the simulation. In Fig. (13), the chromaticity introduces an extra tune shift $\xi_{u}\delta$ to each particle, the same as seen in the left top subgraph in Fig. (8). In Fig. (14), for the result without compensation, the beam loss is rapid, causing the statistical $\varepsilon_{rms,x}$ to decrease. Conversely, the $\varepsilon_{y}$ grows rapidly, resulting in an increase in statistical $\varepsilon_{rms,y}$ despite the beam loss. For the result with raw RDTs, emittance growth could be observed in both horizontal and vertical planes, indicating that the compensation scheme with raw RDTs is not effective under space-charge-induced resonance crossing. For the result with modified RDTs, the $\varepsilon_{rms,x}$ is almost the same as that without errors, and the $\varepsilon_{rms,y}$ grows by 2.5%, which is 0.5% larger than growth observed in Fig. (10). In Fig. (15), for the phase space with modified RDTs, the amplitude of the black halo particle becomes even smaller than that without errors. Fig. (17) demonstrates the $J_{u}$ evolution of this particle, showing that there is no action transfer from the vertical to the horizontal plane. This phenomenon is interesting as the mechanism behind this action decrease under the combined effect of space charge, chromaticity, sextupole errors and corresponding compensation is unclear and warrants further study.
@@ IV CONCLUSIONS
In this paper, the key idea is to propose the space-charge-Twiss modification in RDTs, called modified RDTs. The modification is an approximate approach to obtain the Twiss of particles under the space-charge-induced resonance crossing, through which the modified RDTs could be obtained. The modified RDTs aim to compensate for the combined effect of space charge and magnetic field imperfections, focusing on particles under resonance crossing. This approach is precise for core particles but imprecise for halo particles.
The feasibility of half-integer and 3rd-order resonance crossing compensated by modified RDTs is confirmed by the simulation, in which the raw RDTs and moved RDTs are also demonstrated for comparison. Among these schemes, the modified RDTs exhibits the best performance. The simulations are conducted both with and without chromaticity. The chromaticity has little influence on the overall compensation with modified RDTs, as evidenced by the almost unchanged RMS emittance evolution.